AMS-L^{A}TEX Reference Card #1

See the TEX Reference Card for additional commands. Required packages are indicated as (package).

Document Structure

• Preamble

\documentclass[option(s)]{class}

\usepackage[option(s)]{package(s)}

\begin{document}

• Body

• Front Matter (\frontmatter in book classes)

• Top Matter

\title{. . . }

\title[running head]{… } alternative headline

\date{. . . }

\date{\today} gives current date

\author{. . . }
\maketitle (not in book classes)

• Additional items — ams classes only

\translator{. . . }
\dedicatory{. . . }

\address[optional name]{… }

\curraddress{. . . }

\email[optional name]{… }

\thanks{. . . }
\subjclass{Primary:XXX; Secondary:XXX}
\keywords{. . . }
\thanks{. . . }
\tableofcontents
\chapter{Introduction} (in book classes)

•Abstract (not in book classes)

\begin{abstract}. . . \end{abstract}

• Main Matter (\mainmatter in book classes)

\chapter{. . . }
\section{. . . }
\subsection{. . . }
\appendix

• Back Matter (\backmatter in book classes)

\begin{thebibliography}{99}. . . \end{. . . }
\end{document}

Page Style

\pagestyle{style} set page style to one of:
plain  empty header, page number in footer
empty  empty header and footer
headings  header filled by doc class, empty footer
myheadings  empty footer, fill header with info in

\markboth{lefthead}{righthead}

and \markright{righthead}
\thispagestyle{style} set \pagestyle, only current page
\enlargethispage{\baselineskip}  force an extra line
\renewcommand{\baselinestretch}{2}  doublespaced

fancyheadings package allows custom headers and footers • Page Style Parameters \hoffset, \voffset move page right, down

\paperwidth, \paperheight, \textheight, \textwidth
\topmargin, \headheight, \headsep, \footskip

\pagenumbering{… } e.g., arabic, roman

Classes and Packages

\documentclass[option(s)]{class}

\usepackage[option(s)]{package(s)}

\NeedsTeXFormat{LaTeX2e}[1994/12/01]

•Document Classes

article, book, letter, report, slides
amsart, amsbook, amsproc (all autoload amsmath)

•Useful Packages

amsmath,amsthm,amscd,amssymb,latexsym

fancyheadings allows custom headers and footers alltt all teletype, even \,{,} makeidx,showidx create index, show in margin graphics,graphicx inclusion of graphics enumerate extends the enumerate environment layout shows page layout of doc class multicol flexible multicolumn typesetting showkeys print label keys in margin verbatim extends verbatim environment url typeset URLs allowing line breaks

graphpap  \graphpaper command for \picture environ.

•Document and Package Options

Font Size

8pt, 9pt, 10pt, 11pt, 12pt

Paper Size

a4paper,a5paper,b5paper,legalpaper,letterpaper

Document Preparation

draft,final,notitlepage,titlepage

Page Formatting

onecolumn,twocolumn,oneside,twoside,openany,openright

Equation Numbering

fleqn,leqno,reqno,centertags,tbtags

Equation Limits

intlimits,sumlimits,nonamelimits

AMS (Postscript) Fonts

psamsfonts,noamsfonts

Bibliography (see also BIBTEX)

\begin{thebibliography}{99}. . . \end{. . . }

bibliography with widest label specified

\bibitem{name}  named bibliography item

\bibitem[label]{name} with alternative label to print

\bysame  use long line for same author

\renewcommand{\bibname}{title}  use custom title

\cite{name}  print number of named bib item
\cite[text]{name}  with extra text

Cross Referencing and Numbering

\label{name}  assign label name to numbered item
\ref{name}  print number of named item
\eqref{name}  print number in parentheses (amsmath)
\pageref{name}  print page location of named item
\cite{name}  print number of named bibliography item
\cite[text]{name}  with extra text

\numberwithinsection{equation}{section} number by section

Sectioning and Table of Contents

• Sectioning commands

\command{title}  sectioning command with title
\command[head]{title}  with alternative running head
\command*{title}  with number supressed

    \part       \section          \paragraph
    \chapter    \subsection       \subparagraph

                \subsubsection

\appendix  start appendix

• Table of Contents

\tableofcontents  create and print contents

filename.toc  contents associated to filename.tex

\addcontentsline{toc}{section}{line to add}
\addtocontents{toc}{material to add}
\setcounter{tocdepth}{. . . } set amount to print

Tables and Figures

\begin{table} . . . \caption{text} \label{name} \end{table}

\listoftables  create and print list of tables

\begin{figure} . . . \caption{text} \label{name} \end{figure}

\includegraphics{filename} include image (graphics)

\scaledbox{.5}{\includegraphics{filename}}  scaled graphic

\listoffigures  create and print list of figures

Lists

\item  item within list
\item[label]  item with label

\begin{enumerate}. . . \end{. . . }  numbered items
\begin{itemize}. . . \end{. . . }  bulleted items
\begin{description}. . . \end{. . . } captioned items
\setlength{\itemsep}{0pt}  move items closer

enumerate package  extends enumerate

Displayed Text Material

\begin{center}… \end{… }  centered matrial

\begin{flushright}. . . \end{. . . } flush right matrial
\begin{flushleft}. . . \end{. . . } flush left matrial
\begin{quote}. . . \end{. . . }  short quote
\begin{quotation}. . . \end{. . . } long quote
\begin{verse}. . . \end{. . . }  poetry
\begin{verbatim}. . . \end{. . . }  verbatim material

\verb|… |  verbatim material

\verb*|… | verbatim with spaces marked

verbatim package  extends verbatim

Footnotes, Comments, Other Stuff

\footnote{text}  numbered footnote
%  comment out a line

\begin{comment}… \end{… } long comment (verbatim)

\typeout{text}  print to terminal
\typein{text}  get input from keyboard

\typein[\cmd]{text}  assign input to \cmd

\protect  protects fragile commands
\-  optional hyphen

\hyphenation{hypenated words} extra hyphenated words

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2007 J.H. Silverman, January 2007 v2.0 Math. Dept., Brown Univ., Providence, RI 02912 USA

Permission is granted for noncommercial distribution provided the copyright notice and this permission notice are preserved on all copies.

Dimensions, Spacing, and Glue

Dimensions are specified as ⟨number⟩⟨unit of measure⟩. Glue is specified as ⟨dimen⟩ plus⟨dimen⟩ minus⟨dimen⟩.

point pt pica pc inch in centimeter cm
m width em x height ex math unit mu millimeter mm
1 pc = 12 pt 1 in = 72.72 pt 2.54 cm = 1 in 18 mu = 1 em
\  \quad \qquad  white space (1 space, 1 em, 2 em)
\hspace{10pt}  specified horizontal space
\hspace*{10pt}  space even at line start
Horizontal Spacing (Math): \, thin space \: med space

\; thick space \! neg. thin space \mspace⟨muglue⟩

\strut,\mathstrut  invisible vertical space
\phantom{… }  invisible space
\vphantom{… }  invisible vertical space
\smash[bt]{… }  typeset w/zero height,depth
\hfill  fill with space
\dotfill  fill with dots
\hrulefill  fill with rule (line)
\par  new paragraph
\newline or \\  force a new line
\\*  new line, prohibit page break
\\[5pt]  new line skipping 5 pts
\vspace{1in}  specified vertical space
\vspace*{1in}  space even at page start
\newpage  force a new page

• Length Variables

\newlength{\lngth} create length varible \lngth

\setlength{\lngth}{dimen}  set value of \lngth
\addtolength{\lngth}{dimen} increase \lngth

• Useful Length Assignments

\enlargethispage{\baselineskip} force extra line
\setlength{\hangindent}{30pt}  indentation
\setlength{\hangafter}{3}  indent after

\renewcommand{\baselinestretch}{2} doublespaced

Accents

Type  Example  In Math  In Text

hat  â  \hat              \^

expanding hat  _{abc}  \widehat  none

check  ǎ  \check            \v
tilde  ã  \tilde            \~

expanding tilde  _{abc}  \widetilde  none

acute  á  \acute            \'
grave  à  \grave            \`
dot  ȧ  \dot              \.

double dot  ä  \ddot             \"

breve  ă  \breve            \u
bar  ā  \bar              \=

vector  a⃗  \vec  none
cedilla  ç  none  \c

Additional Text Symbols

\dag †  \copyright         \begin{matrix}c \\ ◯ \end{matrix}  \pounds £
\ddag ‡  \textcircled\{r\}    \begin{matrix}r \\ ◯ \end{matrix}
\P  ¶  \textvisiblespace
\S  §  \textbullet  •

Fonts

• Text Fonts

                                                             \textnormal{. . . } {\normalfont. . . } document font
                                                             \textrm{. . . }     {\rmfamily. . . } roman
                                                             \textsf{. . . }     {\sffamily. . . } sans serif font
                                                             \texttt{. . . }     {\ttfamily. . . }    typewriter style
                                                             \textbf{. . . }     {\bfseries. . . } bold
                                                             \textup{. . . }     {\upshape. . . }  upright
                                                             \textit{. . . }     {\itshape. . . }  italic
                                                             \textsl{. . . }     {\slshape. . . }  slanted
                                                             \textsc{. . . }     {\scshape. . . }  SMALL CAPITALS

\emph{… }       {\em… }  emphasize
\fbox{… }                         framed text

• Font Environments exist for above types, e.g.,

                                                             \begin{ttfamily}. . . \end{. . . }

• Changing Font Sizes

                                                                   \tiny, \scriptsize, \footnotesize, \small
                                                                \normalsize \large, \Large, \LARGE, \huge, \Huge

• Math Fonts

                                                             \mathrm{. . . }  roman

\mathbf{… }  bold (letters)
\boldsymbol{… }  bold (symbol) (amsmath)

                                                             \mathit{. . . }  italic

\mathcal{… }  caligraphic A, B, C

                                                             \usepacakge{eucal} redef \mathcal to script 𝒜, ℬ, 𝒞

\mathfrak{… }  Fraktur 𝔄, 𝔞, 𝔅, 𝔟 (amsfonts)
\mathbb{… }  Blackboard bold 𝔸, 𝔹, ℂ (amsfonts)
\boxed{… }           framed math

• Math Font Sizes

                                                             \displaystyle  display size
                                                             \textstyle  text size

\scriptsize  sub/superscript size
\scriptscriptsize doubly sub/superscripted size

Boxes

\mbox{… }  one line of text
\text{… }  one line of text (amsmath)
\parbox{width}{text}  paragraph of text

\parbox[align][height][inner align]{width}{text}

\marginpar{… }  marginal comment

                                                             \rule[-1pt]{20pt}{10pt}  solid box
                                                             \raisebox{5pt}{text}  raised box

\makebox[width][alignment]{text} box of text
\framebox[width][alignment]{text}    framed text

                                                             \setlength{\fboxsep}{5pt} space around text
                                                             \setlength{\fboxrule}{3pt} width of box borders

Overfull and Underfull Boxes

draft  document class marks overfulls
\overfullrule  width of overfull marker

                                                             \begin{setlength}{\hfuzz}{2pt}. . . \end{. . . }

allow small overfulls

Multicolumn Printing

\twocolumn  double column on new page
\onecolumn  single column on new page

                                                             \begin{multicols}{n}[title]. . . \end{. . . }

multicolumn environment (multicol)

Array and Tabular Environments

                                                                                                                        \begin{tabular}[POS]{COLS}. . . \end{. . . }
                                                                                                                        \begin{array}[POS]{COLS}. . . \end{. . . }

Use tabular for text, array for mathematics

&, \\  column and row separators

POS aligns top (t), bottom (b), center (default)

COLS gives formats for columns:

l,c,r  left, center, right justified
|  vertical rule
@{… }  material between columns
@{}  no space between columns
*{n}{… } n copies of material
p{width} set column width
\hline  horizontal line between rows
\cline{i-j} line across columns i to j

                                                                                                                        \multicolumn{n}{COLS}{. . . }

span n columns using format in COLS

                                                                                                                        \setlength{\tabcolsep}{0pt} set column separation
                                                                                                                        \setlength{\itemsep}{0pt} set item separation
                                                                                                                        \renewcommand{\arraystretch}{1.25} open up array

• Example of a table using \tabular

                                                                                                                        \begin{table}
                                                                                                                        \begin{center}
                                                                                                                        \begin{tabular}{|l|c|c|} \hline
                                                                                                                        Name & Exam & Grade \\ \hline
                                                                                                                        Dan & 97\% & A \\ \hline
                                                                                                                        \end{tabular}
                                                                                                                        \caption{Math 101 Final Grades}
                                                                                                                        \label{GradeTable}
                                                                                                                        \end{center}
                                                                                                                        \end{table}

Name Exam Grade
Dan  97%  A

Math 101 Final Grades

Tabbing Environment

\begin{tabbing}… \end{… } tabbing environment
\=  set tab
\\  end line
\>  move to next tab
\kill  do not print line

File Suffixes and Types

• L^{A}TEXSource Files

.tex  File containing a L^{A}TEXdocument

.sty, .cls L^{A}TEX style and document class files

.fd  Font definition file

• Files Written by L^{A}TEX (See also BIBTEXand MAKEINDEX)

.aux  cross-referencing and list information
.dvi  device independent typeset file
.glo  list of glossary entries
.lof  list of figures (read by \listoffigures)
.lot  list of tables (read by \listoftables)
.toc  table of contents (read by \tableofcontents)
.log  L^{A}TEX log file

\nofiles supresses all except .log and .dvi

\begin{matrix}c \\ ◯ \end{matrix} 2007 J.H. Silverman, Permissions on back. v2.0

Send comments and corrections to J.H. Silverman, Math. Dept., Brown Univ., Providence, RI 02912 USA. ⟨jhs@math.brown.edu⟩

AMS-L^{A}TEX Reference Card #2

See the TEX Reference Card for additional commands. The notation (package) indicates a required package.

Math Environments

\(… \) or $… $  inline math
\[… \] or $$… $$  displayed math

\begin{equation}\label{eqname}. . . \end{. . . }

numbered and labeled equation

\ref{eqname}  refer to labeled eqn
\mbox{… }  text in math

• The following require amsmath

\text{… }  text in math

\begin{equation*}. . . \end{. . . } unnumbered eqn

\tag{eqtag} use eqtag instead of number

\notag  supress equation tag
\eqref{eqname}  ref with parens

\begin{subequations}. . . \end{. . . }

group equations for numbering

\numberwithin{equation}{section}

number equations within sections

Theorems, Lemmas, Etc.

• Defining Theorem-Like Environments

\newtheorem{name}{label} theorem environment
\newtheorem*{name}{label} unnumbered (amsthm)

\newtheorem{name}[other name]{label}

numbered consecutively with other environment

\newtheorem{name}{label}[section]

numbered by section (or chapter, etc.)

\swapnumbers  put numbers on left

• Theorem-Like Environment Styles (amsthm)

\theoremstyle{plain}  most emphatic
\theoremstyle{defintion} medium emphasis
\theoremstyle{remark}  least emphatic

• Invoking Theorem-Like Environments

\begin{name}… \end{… } invoke environment
\begin{name}[label]…  invoke with new label

If proclamation starts with a list, put in \hfill

\begin{proof}… \end{… } proof environment

\begin{proof}[label]… \end{… } proof with label

\qedsymbol  end of proof marker

\renewcommand{\qedsymbol}{. . . }  redefine marker

Commutative Diagrams (amscd)

Separate lines with \\, do not use &s \begin{CD}… \end{CD} commutative diagram

@>#1>#2>  right arrow with labels
@<#1<#2<  left arrow with labels
@V#1V#2V  down arrow with labels
@A#1A#2A  up arrow with labels
@=  long horizontal equal sign
@|  long vertical equal sign
@.  leave out an arrow

Multiline Math Displays (amsmath)

Use as \begin{command}. . . \end{command}
Separate items with &, separate lines with \\
No \\ on last line, \\[dim] to skip space

• Full Math Environments (full line)

gather  centered, numbered equations
gather*  centered, unnumbered equations
multline first line left, last line right, rest centered
multline* same as multline, but unnumbered
align  formulas aligned at & signs
align*  same as align, but unnumbered
flalign  flush left and right align
alignat    align without space, needs

             argument \begin{alignat}{# of cols}

\intertext{text}  text between lines

\shoveleft,\shoveright move multline line left, right

\allowdisplaybreaks  allow page breaks (\\* prohibits)
\displaybreak  force page break (before \\)

• Math Subenvironments (within math display)

gathered centered equations
aligned formulas aligned at & signs
split  split long formula within other environment
cases  cases, with { on left
matrix  matrix (of up to 10 columns)

pmatrix, bmatrix, vmatrix, Vmatrix

matrix variants enclosed by (⋯), [⋯], | ⋯ |, ∥ ⋯ ∥

\setcounter{MaxMatrixCols}{12}

increase number of matrix columns

\hdotsfor{num} dots across columns

Overlines, Underlines, and Arrows

\underline{. . . }  underline
\overline{. . . }  overline
\overbrace{. . . }^{. . . }  overbrace
\underbrace{. . . }_{. . . }  underbrace

\overightarrow{… }  over right arrow

\overleftarrow{. . . }  over left arrow

\overleftrightarrow{… } over left-right arrow

\underrightarrow{. . . },\underleftarrow{. . . }, etc.

\xrightarrow[bot]{top}  stretchable w/sub/supscripts
\xleftarrow[bot]{top}  stretchable w/sub/supscripts

Operator Names

\arccos  \cos   \csc  \exp  \ker  \liminf  \min  \sinh
\arcsin  \cosh  \deg  \gcd  \lg   \limsup  \Pr   \sup
\arctan  \cot   \det  \hom  \lim  \log     \sec  \tan
\arg     \coth  \dim  \inf  \ln   \max     \sin  \tanh
a \equiv b \pmod{m} a ≡ b (mod m)
a \equiv b \mod{m} a ≡ b mod m
a \bmod m  a mod m
\DeclareMathOperator{\cmd}{opname} create operator
\DeclareMathOperator*{\cmd}{opname}  with limits

\operatorname{… } typeset as an operator

\operatorname*{. . . }  with limits

Large Operators

                                                             ∑ \sum  ⋂ \bigcap  ⨀ \bigodot
                                                             ∏ \prod  ⋃ \bigcup  ⨂ \bigotimes
                                                             ∐ \coprod  ⨆ \bigsqcup  ⨁ \bigoplus
                                                             ∫ \int  ⋁ \bigvee  ⨄ \biguplus
                                                             ∮ \oint  ⋀ \bigwedge

\substack{xxx\\ yyy}  stacked sub or superscripts
\limits,\nolimits  force or forbid displayed limits

                                                             \oint,\iint,\iiint,\iiiint,\idotsint

integral variants (amsmath)

Delimiters

                                                             [ \lbrack or \[  { \lbrace or \{  ⟨ \langle
                                                             ] \rbrack or \]  } \rbrace or \}  ⟩ \rangle
                                                             | \vert or \|  ⌊ \lfloor  ⌈ \lceil
                                                             ∥ \Vert or \|  ⌋ \rfloor  ⌉ \rceil
                                                             ↑ \uparrow  ⇑ \Uparrow  ↕ \updownarrow
                                                             ↓ \downarrow  ⇓ \Downarrow  ⇕ \Updownarrow

\left(  \right)  expanding delimiters
\left.  \right.  empty delimiters
\bigl(  \bigr)  big delimiters
\Bigl(  \Bigr)  bigger delimiters
\biggl(  \biggr)  even bigger delimiters
\bigm|,\biggm|  big binary relation delimiters

Roots

\sqrt{… }  square root √
\sqrt[n]{… }  nth root ^{^{n} √}

                                                             \leftroot{2},\uproot{2} move root left or up

Ellipses

                                                             \ldots,\cdots,\dots ellipses

\vdots,\ddots  vertical and diagonal dots

                                                             \dotsc,\dotsb,\dotsm,\dotsi more ellipses (amsmath)

Fractions and Stacked Relations

\frac\{n\}\{d\}  fraction \frac{n}{d}
\dfrac{n}{d}  displaystyle fraction
\tfrac{n}{d}  textstyle fraction
\binom\{n\}\{d\}  binomial coefficient \binom{n}{d}

\genfrac{ldelim}{rdelim}{thick}{style}{num}{den}

\cfrac{… }{… }  continued fraction
\stackrel{top}{bot} stacked relation
\overset{top}{bot} stacked symbol (amsmath)
\underset{bot}{top} stacked relation (amsmath)

                                                             \sideset{_{ll}^{ul}}{_{lr}^{ur}}{largeop}

large operator with left/right sub/supscripts

Negated Relations

\not  negate a relation
\ne  not equal ≠
\notin not a member of ∉
\nmid not divisible ∤

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2007 J.H. Silverman, January 2007 v2.0 Math. Dept., Brown Univ., Providence, RI 02912 USA Permission is granted for noncommercial distribution provided the copyright notice and this permission notice are preserved on all copies.

User Defined Commands

\newcommand{\name}{replacement text} new command

\newcommand{\name}[n]{text with #1,#2,. . .,#n}

new command with n arguments

Example: \newcommand{\vect}[2]{#1_1,\ldots,#1_{#2}}
\newcommand{\name}[n][default]{. . . }

command with args and default value for #1

\renewcommand{… }{… }  redefine existing command
\providecommand{… }{… } define if doesn’t exist
\newcommand*{… }{… }  command with one par arg
\ensuremath{… }  forces math mode

\show\command  print definition of \command

\showthe\paramname  print value of a parameter

User Defined Environments

\newenvironment{name}{pretext}{posttext}

new environment with material before and after

\newenvironment[n]{name}{. . . }{. . . }

environment with n arguments

\newenvironment[n][default]{name}{. . . }{. . . }

environment with default value for #1

\renewenvironment{name}{… }{… } redefine envrment

MAKEINDEX

• MakeIndex File Suffixes

.idx, .ind, .ilg  entry listing, index file, log file

• MakeIndex Commands in Document File

\usepackage{makeidx}  use indexing package

(Do not include this line if using AMS packages.)

\makeindex tell L^{A}TEX to create an .idx file
\printindex tell L^{A}TEX to print index here
\nofiles  supresses creation of .idx and .glo files

• Creating MakeIndex .idx File

\index{entry}  main entry
\index{entry!entry}  subentry
\index{entry!entry!entry}  subsubentry
\index{text@entry}  with placement info
\index{entry|see{entry}}  cross referenced entry

\index{entry|modifier} entry with page modifier

e.g. \index{gnats|textbf} give bold page number \index{entry|(} … \index{entry|)} page range Special Characters: "! "@ "| ""

• Creating An Index With MakeIndex

(1) Typeset document containing \makeindex command.

(2) Run MakeIndex on .idx file to create .ind file.

(3) Typeset document containing \printindex command.

Glossary

\makeglossary  tell L^{A}TEX to create a .glo file
\glossary{entry}  create a glossary entry

\glossaryentry{entry}{page no.} entries in .glo file

\input filename.glo read glossary file

User must define \makeglossary, e.g.,

\newcommand{\glossaryentry}[2]{#1, page #2\par}

Time and Date

\today current date

Use \the to display the following items \day, \month, \year, \time (minutes since midnight)

Counters

\newcounter{cntr}  create new counter named cntr
\newcounter{cntr}[cntr1]reset cntr when cntr1 changes

\setcounter{cntr}{value} set value of cntr
\stepcounter{cntr}  increment cntr
\refstepcounter{cntr}  increment and reset \label
\addtocounter{cntr}{n} increment by n
\value{cntr}  value stored in \cntr
\thecntr  the value of cntr

calc package to do counter arithmetic • Counter Styles

\arabic{} \roman{} \Roman{} \alph{} \Alph{}

• Standard Counters

equation footnote figure page table
part chapter section subsection subsubsection
paragraph subparagraph enumi enumii enumiii enumiv

secnumdepth  depth to which sections are numbered
tocdepth  depth to which sections are put into toc

Customized List Environments

\begin{list}{default label}{declarations} \item item 1 text \item item 2 text

\end{list}
\begin{trivlist}. . . \end{trivlist}

list with no labels or declarations, trivial lengths

•Declarations \setlength{length parameter}{length}

\usecounter{counter name}

[Create counter first using \newcounter{counter name}.] •Length Parameters (see page 113 of Lamport for more)

\topsep  separate preceding text and first item
\itemsep  separate items
\leftmargin indent of item box from left margin
\labelwidth width of box for item label
\labelsep  separate label box from item box

The picture Environment

\begin{picture}(w,h). . . \end{picture} picture
\begin{picture}(w,h)(Δx,Δy). . .  with offset

\put(x,y){picture object}  place object

\multiput(x,y)(Δx,Δy){n}{object}  n times

Picture Objects:

\makebox(x,y)[tblr]{text}  box with text

\line(Δx,Δy){x length}  line of slope Δy/Δx
\vector(Δx,Δy){x length}  arrow of slope Δy/Δx
\circle{r}  circle of radius r
\circle*{r}  filled circle
\oval(x,y)[lrtb]  oval (part or whole)

\shortstack{abc\\xyz\\}  stacked text
\framebox(x,y)[tblr]{text}  framed text

\frame{text},fbox{text}  other framed boxes
\dashbox{d}(x,y){text}  dashed box
\qbezier(x_{1},y_{1})(x_{2},y_{2})(x_{3},y_{3}) quadratic curve

\savebox{\name}(x,y){. . . }  store material

\usebox{\name}  retrieve material
\graphpaper[n]{x,y}{w,h}  print grid (graphpap)

\setlength{\unitlength}{1pt} change size of picture
\thinlines,\thicklines  adjust line thickness

Color (color)

\color{color}  change color
\textcolor{color}{text}  colored text
\colorbox{color}{text}  colored background
\fcolorbox{col_{1}}{col_{2}}{text}  colored border & background

\setlength{\fboxsep}{5pt}  put space around text
\setlength{\fboxrule}{3pt}  width of border of box

\pagecolor{color}  set background color of page
\definecolor{name}{rgb}{r, g, b} define an RGB color

\definecolor{name}{cmyk}{c, m, y, k} define a CMYK color Predefined Colors

  black, white, red, green, blue, yellow, cyan, magenta

BIBTEX

• BIBTEX File Suffixes

.bib  BIBTEX bibliographic database file
.bst  BIBTEX bibliographic style file
.blg  BIBTEX log file
.bbl  BIBTEX document bibliography file

• BIBTEX Commands in Document File

\bibliographystyle{bib style file}
  Examples: plain, amsplain, unsrt, alpha, abbrv

\bibliography{bib database file(s)}

\cite{label}  cite a reference
\nocite{label} include ref in bib without citation
\nocite{*}  include all references in bibliography

• Creating BIBTEX Database File @STRING{name = "text"} define an abbreviation Put braces around non-initial capitalized title words. Use and to separate multiple authors in author field

•General Format of a Database Entry

@ENTRYTYPE{label,
    fieldtype1 = {entry1},
    fieldtype2 = {entry2},

⋮ }

•Database Entry Types

@ARTICLE{. . . }         @MASTERSTHESIS{. . . }
@BOOK{. . . }            @MISC{. . . }
@BOOKLET{. . . }         @PHDTHESIS{. . . }
@INBOOK{. . . }          @PROCEEDINGS{. . . }
@INCOLLECTION{. . . }    @TECHREPORT{. . . }
@INPROCEEDINGS{. . . }   @UNPUBLISHED{. . . }
@MANUAL{. . . }          @COMMENT{. . . }

•Field Types Within Entries

address
author
booktitle
chapter
crossref
edition

           editor
           howpublished
           institution
           journal
           key
           language

                         month
                         note
                         number
                         organization
                         pages
                         publisher

                                      school
                                      series
                                      title
                                      type
                                      volume
                                      year

• Creating Document Bibliography With BIBTEX

(1) Typeset document to get new .aux file.

(2) Run BIBTEX on .aux file to create .bbl file.

(3) Retypeset document twice.

\begin{matrix}c \\ ◯ \end{matrix} 2007 J.H. Silverman, Permissions on back. v2.0

Send comments and corrections to J.H. Silverman, Math. Dept., Brown Univ., Providence, RI 02912 USA. ⟨jhs@math.brown.edu⟩